OpenStudy (anonymous):

If g(x) = x2 − 1, find g(3)

jimthompson5910 (jim_thompson5910):

Replace x with 3 and evaluate.

jimthompson5910 (jim_thompson5910):

g(x) = x^2 - 1 g(3) = 3^2 - 1

OpenStudy (anonymous):

2 is actually an exponent

jimthompson5910 (jim_thompson5910):

I know

jimthompson5910 (jim_thompson5910):

that's why I wrote x^2, the ^ means the stuff after it is an exponent

OpenStudy (anonymous):

oh okay lol thank you

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

so g(3) = 8?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

you're very welcome